# Hallmarks of Optimization Trajectories in Neural Networks and LLMs: The Lengths, Bends, and Dead Ends

### v1

@@ 6 Related Work and Discussion
Prior work has theoretically noted a notion of directional convergence (Ji & Telgarsky, 2020), wherein the parameters of simple networks and classifiers converge quickly, in terms of their direction. Likewise (Merrill et al., 2020) have observed that cosine similarity between subsequent parameter checkpoints during T5 (Raffel et al., 2023) pre-training rapidly approaches $1$. Our analysis of the trajectory maps can thus be seen as related to this aspect, however, we carry out a comprehensive analysis of how directional similarity of the parameters evolves over the course of training, how it behaves in a variety of relevant settings triggered by enabling or disabling key hyperparameters, and how it varies across models of unprecedented scales. Moreover, our analysis of the trajectory via the quantitative and qualitative hallmarks provides a more refined and insightful picture of the nature of optimization trajectory, instead of the mere fact about the rapid directional convergence noted in prior works.
@@
Implicit bias (Gunasekar et al., 2018; Li et al., 2019, 2020; Moroshko et al., 2020) has emerged as one of the main contenders for explaining the success of deep neural networks. This principle, or ideology, has also inspired several works which seek to uncover the implicit effects that might be latent in the regular working of hyperparameters. To name a few, Andriushchenko et al. (2023) for instance suggest a loss stabilization mechanism behind weight decay, while Liu et al. (2023) attempt to characterize the implicit bias of large learning rates in terms of resulting in a flatter solution, and Jelassi & Li (2022); Cao et al. (2023) explain the implicit bias of momentum and batch normalization with regards to margin. In contrast, we take a broad and general perspective by adopting a full trajectory view. In particular, we explore more deeply an intriguing interaction of various optimization knobs, such as that between momentum and weight decay, and show its link with Edge of Stability (Cohen et al., 2022).
Complexity of Optimization Path as a Generalization Measure?
There has been, of course, a huge set of works analyzing the generalization behaviour but with more of a focus on final basin quality measures, like sharpness (Zhang et al., 2021; Petzka et al., 2021; Singh et al., 2022; Gastpar et al., 2023). While apriori, we might believe that the final basin quality matters more than the particular path taken to reach it (however complex the path might be). But, thinking about the implicit bias (whatever that truly is), which is at play when optimizing over-parameterized networks, and considering our empirical findings about the trajectory maps when the hyperparameters are tweaked, it seems that the kind of optimization path followed might matter more than the eventual destination. In fact, our notion of the mean directional score (MDS) has shown promise to be a useful measure to track the quality of optimization, given its well-behaved nature in various controlled settings. While, here our primary focus was to reveal, much more broadly, the varied characteristics of the optimization trajectory, it would form, without a doubt, a fascinating question to explore MDS as an alternative correlator/predictor of generalization in future work.
@@
Besides, our analysis can also pave the way for a novel data-free mechanistic understanding of LLMs (see for instance, Figure 9, where the query-key-value weights for the middle layer interestingly converge the latest; Appendix E contains the entire set of figures.). This holistic optimization path perspective can complement the top-down approach of mechanistic understanding of LLMs via influence functions (Grosse et al., 2023) and the bottom-up circuit view of Transformers (Elhage et al., 2021).
Exploiting Redundancy in Optimization.
Lastly, given the high cosine similarity between the parameter checkpoints, this raises the exciting prospect of exploiting the structure of optimization trajectories for faster optimization, like via a line search (Vaswani et al., 2019) or adapting only scalar parameters per layer (Frankle et al., 2021) such as those in LayerNorm. Further, by carrying out a layerwise directional analysis, we might be able to ‘pause’ optimization at the level of certain layers or blocks of parameters a little while into training, or simply, intersperse their updates at some intervals. All in all, this would allow for a more structured form of efficient low-rank training (Hu et al., 2021) in otherwise extremely high dimensional spaces.
@@ 7 Conclusion
In sum, we present and develop the perspective of studying the entire optimization trajectories, and their directional geometry, for analyzing the nature of solutions reached by neural networks (including LLMs) in different settings. More specifically, we consider qualitative and quantitative hallmarks (such as the trajectory map and angle between successive trajectory points, respectively), and utilize them to showcase the manner in which key hyperparameters and regularization techniques as well as the network scale affect the ensued optimization trajectories. Notably, we observe that increasing momentum, weight decay, batch size, and learning rate increase the extent of directional exploration in the trajectory, while sharpness-aware minimization and increased network scale have an implicit regularization effect of reducing the extent of directional exploration.
Nevertheless, we have merely scratched the surface of this trajectory perspective into understanding optimization and generalization behaviour in neural networks. We genuinely believe that there is a lot to be understood about the complex, intermingled behaviour of optimization in deep learning, and hopefully, this work will bring further nuance in these areas and foster an understanding of the structure of these trajectories.
Impact Statement.
If the directional redundancy present in the optimization paths of neural networks and large language models can be tapped, this would enable faster algorithmic techniques of training large networks quite efficiently and thereby help towards reducing the otherwise heavy carbon footprint of these models. Besides, the hallmarks of optimization trajectories that we presented can help towards furthering our understanding of the generalization aspects of these networks.
@@ Appendix A Omitted Proofs
Given a quadratic problem with $\ell_{2}$ regularization of strength $\alpha>0$, namely,
$\min_{{\boldsymbol{\theta}}\in{\mathbb{R}^{d}}}\frac{1}{2}{\boldsymbol{\theta}}^{\top}{\bf M}{\boldsymbol{\theta}}+\frac{1}{2}\alpha\|{\boldsymbol{\theta}}\|^{2}\,,$
with ${\bf M}$ symmetric, the angle between successive steps $\Delta_{t}={\boldsymbol{\theta}}_{t}-{\boldsymbol{\theta}}_{t-1},\Delta_{t+1}={\boldsymbol{\theta}}_{t+1}-{\boldsymbol{\theta}}_{t}$, when using gradient descent with a one-step momentum ($\mu>0$) and learning rates $\eta_{t},\eta_{t+1}$, can be upper and lower bounded as follows:
$\langle\Delta_{t},\Delta_{t+1}\rangle\leq\eta_{t}\eta_{t+1}(1-\mu\eta_{t}-\eta_{t}\alpha-\eta_{t}\lambda_{d})(\lambda_{d}+\alpha)^{2}\|{\boldsymbol{\theta}}_{t-1}\|^{2}$
$\langle\Delta_{t},\Delta_{t+1}\rangle\geq\eta_{t}\eta_{t+1}(1-\mu\eta_{t}-\eta_{t}\alpha-\eta_{t}\lambda_{1})(\lambda_{1}+\alpha)^{2}\|{\boldsymbol{\theta}}_{t-1}\|^{2}$
@@
where, we have denoted the eigenvalues of ${\bf M}$ as $\lambda_{1}\geq\cdots\geq\lambda_{d}$. Since the inner product of the updates is a quadratic form, we can upper and lower bound it based on the maximum and minimum eigenvalues of ${\bf Z}$, thus giving:
@@
Because of the above form of eigenvalues of ${\bf Z}$ (diagonal matrices have their eigenvalues as their diagonal entries), we will have: $\lambda_{\max}({\bf Z})=(1-\mu\eta_{t}-\eta_{t}\alpha-\eta_{t}\lambda_{d})(\lambda_{d}+\alpha)^{2}$ and $\lambda_{\min}({\bf Z})=(1-\mu\eta_{t}-\eta_{t}\alpha-\eta_{t}\lambda_{1})(\lambda_{1}+\alpha)^{2}$
@@ Appendix B Why cosine similarities increase with scale?
Note, we assume that the majority of the parameter norm lies in the square hidden matrices, and not the input or output layers. Moreover, we use $o,O,\Theta$ to denote standard mathematical notation with regards to scaling in the limit width $n\rightarrow\infty$. For vectors, this notation is entry-wise.
@@
We suppose $x$ has $\Theta(1)$ entries, as is the case with standard initialisations/parameterisations (He et al., 2015). We suppose $W_{0}$ has i.i.d. elements with initialisation that is $O(1/\sqrt{n})$ in order to ensure that each element of the features $h$ has entries $\Theta(1)$.
@@
For our features to be stable (i.e. $\Theta(1)$) after the update, we need $n\eta\text{d}h$ to be $O(1)$, because $\frac{x_{0}^{\top}x_{0}}{n}=\Theta(1)$ by assumption on $x$. NB: if $n\eta\text{d}h=o(1)$ we have no feature learning (ie NTK regime (Jacot et al., 2018)), and if $n\eta\text{d}h=\Theta(1)$ we have feature learning (ie $\mu$P (Yang et al., 2022)).
@@
But because $W_{0}=\Theta(1/\sqrt{n})$, the initialisation will elementwise-dominate the $O(1/n)$ update for the first training step (and more training steps follows by induction). As a result, the update $W_{T}-W_{0}$ will always be an order of at least $\sqrt{n}$ smaller than the initialisation, and hence the new parameters $W_{T}$ will be exactly aligned with the initialisation $W_{0}$ for all $T$ in the large width limit, i.e. the cosine similarities will be $1$.
Appendix C Experimental Results
C.1 ResNet50: Switching off the hyperparameters
@@
Figure 11: Relative Trajectory Maps, with respect to initialization, of ResNet50 models across different amounts of momentum and weight decay
@@
C.2 ResNet50: Weight Decay, SGD
@@
C.3 ResNet50: Weight Decay, AdamW
@@
C.4 ViT: Weight Decay, AdamW
(a) $\angle(\theta_{t+1}-\theta_{t},\theta_{t})$
(b) $\angle(\theta_{t+1}-\theta_{t},\theta_{T}-\theta_{0})$
(c) $\angle(\theta_{t+k}-\theta_{t},\theta_{t}-\theta_{t-k})$, for $k=1$
(d) $\angle(\theta_{t}-\theta_{0},\theta_{T}-\theta_{0})$
(e) $\angle(\theta_{t+1}-\theta_{t},\theta_{T}-\theta_{0})$
(g) Apex Angle at Origin $\angle(\theta_{t},\theta_{0})$
(a) $\|\theta_{t}\|_{2}$
(b) $\|\theta_{t+k}-\theta_{t}\|_{2}$
(c) $\|\theta_{t}-\theta_{0}\|_{2}$
(a) Eigenvalues: ${\bf K}$
(b) Eigenvalues: ${\bf K}_{0}$
(c) Eigenvalues: ${\bf C}$
(d) Eigenvalues: ${\bf C}_{0}$
C.5 ResNet50: Sharpness Aware Minimization analysis
@@
C.6 ResNet50: Momentum Analysis, LR 0.1, WD 0.0001
@@
C.11 VGG16 Batch Size Analysis
Figure 60: Trajectory Maps of VGG16 models across different batch sizes

### v2

@@ 7 Related Work
Prior work has theoretically noted a notion of directional convergence [Ji and Telgarsky, 2020], wherein the parameters of simple networks and classifiers converge quickly, in terms of their direction. Likewise [Merrill et al., 2020] have observed that cosine similarity between subsequent parameter checkpoints during T5 [Raffel et al., 2023] pre-training rapidly approaches one. Our analysis of the trajectory maps can thus be seen as related to this aspect, however, we carry out a comprehensive analysis of how directional similarity of the parameters evolves over the course of training, how it behaves in a variety of relevant settings triggered by enabling or disabling key hyperparameters, and how it varies across models of unprecedented scales. Moreover, our analysis of the trajectory via the quantitative and qualitative hallmarks provides a more refined and insightful picture of the nature of optimization trajectory, instead of the mere fact about the rapid directional convergence noted in prior works.
@@
Implicit bias [Gunasekar et al., 2018; Li et al., 2019, 2020; Moroshko et al., 2020] has emerged as one of the main contenders for explaining the success of deep neural networks. This principle has also inspired several works which seek to uncover the implicit effects that might be latent in the regular working of hyperparameters. To name a few, Andriushchenko et al. [2023] for instance suggest a loss stabilization mechanism behind weight decay, while [Liu et al., 2023] attempt to characterize the implicit bias of large learning rates in terms of resulting in a flatter solution, and Jelassi and Li [2022]; Cao et al. [2023] explain the implicit bias of momentum and batch normalization with regards to margin. In contrast, we take a broad and general perspective by adopting a full trajectory view. In particular, we explore more deeply an intriguing interaction of various optimization knobs, such as that between momentum and weight decay, and show its link with Edge of Stability [Cohen et al., 2022].
@@
Besides, our analysis can also pave the way for a novel data-free mechanistic understanding of LLMs (as shown in Figure 8. This holistic optimization path perspective can complement the top-down approach of mechanistic understanding of LLMs via influence functions [Grosse et al., 2023] and the bottom-up circuit view of Transformers [Elhage et al., 2021].
@@ 8 Conclusion
Overall, we have merely scratched the surface of this trajectory perspective into understanding optimization behaviour in neural networks. We genuinely believe that there is a lot to be understood about the complex, intermingled behaviour of optimization in deep learning, and hopefully, this work will bring further nuance in these areas and contribute towards hybrid optimization schemes that can exploit the showcased directional redundancy.
Limitations and Future Work. While we have focused on the directional aspects of trajectories, the length of the trajectories also holds relevance, especially when the experiments under comparison may have unequal number of sampled checkpoints. We expect a version of MDS, which is weighted by the step lengths to be useful in such a scenario. Besides, our experiments with training scalar parameters are currently based on vision models, given the excessive resources involved in testing this for LLMs. However, we hope that the wider community can seize on this observation and exploit it for training LLMs more efficiently.
Acknowledgments and Disclosure of Funding
We would like to thank Tiago Pimentel for reading a draft of the paper and providing useful comments. Sidak Pal Singh would also like to acknowledge the financial support from Max Planck ETH Center for Learning Systems.
@@ Appendix A Omitted Proofs
Given a quadratic problem with $\ell_{2}$ regularization of strength $\alpha>0$, namely, $\min_{{\boldsymbol{\theta}}\in{\mathbb{R}^{d}}}\frac{1}{2}{\boldsymbol{\theta}}^{\top}{\bf M}\,{\boldsymbol{\theta}}+\frac{1}{2}\alpha\|{\boldsymbol{\theta}}\|^{2}\,,$ with ${\bf M}\in{\mathbb{R}^{d\times d}}$ symmetric with eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{d}$, the angle between successive steps $\Delta_{t}={\boldsymbol{\theta}}_{t}-{\boldsymbol{\theta}}_{t-1},\Delta_{t+1}={\boldsymbol{\theta}}_{t+1}-{\boldsymbol{\theta}}_{t}$, when using gradient descent with a one-step momentum ($\mu>0$) and learning rates $\eta_{t},\eta_{t+1}$, can be upper and lower bounded as follows:
$\langle\Delta_{t},\Delta_{t+1}\rangle\leq\eta_{t}\eta_{t+1}(1-\eta_{t}(\mu+\alpha+\lambda_{d}))(\lambda_{d}+\alpha)^{2}\|{\boldsymbol{\theta}}_{t-1}\|^{2}$
$\langle\Delta_{t},\Delta_{t+1}\rangle\geq\eta_{t}\eta_{t+1}(1-\eta_{t}(\mu+\alpha+\lambda_{1}))(\lambda_{1}+\alpha)^{2}\|{\boldsymbol{\theta}}_{t-1}\|^{2}$
@@
where, we have denoted the eigenvalues of ${\bf M}$ as $\lambda_{1}\geq\cdots\geq\lambda_{d}$.
Since the inner product of the updates is a quadratic form, we can upper and lower bound it based on the maximum and minimum eigenvalues of ${\bf Z}$, thus giving:
@@
Because of the above form of eigenvalues of ${\bf Z}$ (diagonal matrices have their eigenvalues as their diagonal entries), we will have:
$\lambda_{\max}({\bf Z})=(1-\mu\eta_{t}-\eta_{t}\alpha-\eta_{t}\lambda_{d})(\lambda_{d}+\alpha)^{2}$ and $\lambda_{\min}({\bf Z})=(1-\mu\eta_{t}-\eta_{t}\alpha-\eta_{t}\lambda_{1})(\lambda_{1}+\alpha)^{2}$
@@ Appendix B Why cosine similarities increase with scale?
Note, we assume that the majority of the parameter norm lies in the square hidden matrices, and not the input or output layers. Moreover, we use $o,O,{\boldsymbol{\theta}}$ to denote standard mathematical notation with regards to scaling in the limit width $n\rightarrow\infty$. For vectors, this notation is entry-wise.
@@
We suppose $x$ has ${\boldsymbol{\theta}}(1)$ entries, as is the case with standard initialisations/parameterisations He et al. [2015]. We suppose $W_{0}$ has i.i.d. elements with initialisation that is $O(1/\sqrt{n})$ in order to ensure that each element of the features $h$ has entries ${\boldsymbol{\theta}}(1)$.
@@
For our features to be stable (i.e. ${\boldsymbol{\theta}}(1)$) after the update, we need $n\eta\text{d}h$ to be $O(1)$, because $\frac{x_{0}^{\top}x_{0}}{n}={\boldsymbol{\theta}}(1)$ by assumption on $x$. NB: if $n\eta\text{d}h=o(1)$ we have no feature learning (ie NTK regime Jacot et al. [2018]), and if $n\eta\text{d}h={\boldsymbol{\theta}}(1)$ we have feature learning (ie $\mu$P Yang et al. [2022]).
@@
But because $W_{0}={\boldsymbol{\theta}}(1/\sqrt{n})$, the initialisation will elementwise-dominate the $O(1/n)$ update for the first training step (and more training steps follows by induction). As a result, the update $W_{T}-W_{0}$ will always be an order of at least $\sqrt{n}$ smaller than the initialisation, and hence the new parameters $W_{T}$ will be exactly aligned with the initialisation $W_{0}$ for all $T$ in the large width limit, i.e. the cosine similarities will be $1$.
Appendix C Comparing Gradient Trajectories with Random Walks
The structure that we observe in trajectory maps following gradient trajectories raises the question of if we would observe similar structure in a random walk.
If we have $T$ timesteps or epochs, with parameter space ${\boldsymbol{\theta}}\in\mathbb{R}^{p}$ and “learning rate” schedule $(\eta_{t})_{t=1}^{T}$, we can consider a random walk with updates:
[…]
which is to say that at time step $t$, each parameter coordinate in the parameter vector is updated independently with a Gaussian of variance $\eta_{t}^{2}$, and the updates are independent across different time steps.
Then, if $\theta^{i}$ denotes a single parameter coordinate for a dimension $i\leq p$, for two time steps $s<t$, we have:
$(\theta^{i}_{s},\theta^{i}_{t})\sim\mathcal{N}(0,\bigl{(}\begin{smallmatrix}H_{s}&H_{s}\\ H_{s}&H_{t}\end{smallmatrix}\bigr{)})$
where $H_{u}=\sum_{t^{\prime}=1}^{u}\eta_{t^{\prime}}^{2}$ is the cumulative squared learning rate from $t^{\prime}=1$ to $t^{\prime}=u$.
Then, by the strong law of large numbers we have for the large parameter space $p\rightarrow\infty$ limit:
$\frac{1}{p}\lVert{\boldsymbol{\theta}}_{s}\rVert_{2}^{2}=\frac{1}{p}\sum_{i=1}^{p}(\theta^{i}_{s})^{2}\overset{a.s.}{\rightarrow}H_{s}\,,\quad\frac{1}{p}\lVert{\boldsymbol{\theta}}_{t}\rVert_{2}^{2}=\frac{1}{p}\sum_{i=1}^{p}(\theta^{i}_{t})^{2}\overset{a.s.}{\rightarrow}H_{t}$
$\frac{1}{p}\langle{\boldsymbol{\theta}}_{s},{\boldsymbol{\theta}}_{t}\rangle=\frac{1}{p}\sum_{i=1}^{p}\theta^{i}_{t}\theta_{s}^{i}\overset{a.s.}{\rightarrow}H_{s}$
and by the property of composing almost sure limits, we also have almost sure convergence in the cosine similarity:
$\frac{\langle{\boldsymbol{\theta}}_{s},{\boldsymbol{\theta}}_{t}\rangle}{\lVert{\boldsymbol{\theta}}_{s}\rVert_{2}\lVert{\boldsymbol{\theta}}_{t}\rVert_{2}}\overset{a.s.}{\rightarrow}\frac{H_{s}}{\sqrt{H_{s}H_{t}}}=\sqrt{\frac{H_{s}}{H_{t}}}$
in the large parameter space limit, which we use as an approximation to give analytic formulas for the trajectory map and MDS that we can compare to gradient trajectories.
One thing to note, is that this cosine similarity $\sqrt{\frac{H_{s}}{H_{t}}}$ becomes invariant to the scale of the learning rates $\eta$, and instead it is the relative rate of decay in the learning rate schedule that matters.
C.1 Experimental Results
Figure 11: With step size decay: Relative Trajectory Map for a Random Walk/Brownian motion in both analytic and empirically simulated settings.
The first thing to note is that our empirical simulation of the random walk matches the theoretical limit described in the section above, for a finite parameter count (such as $10,000$). Next, comparing the these relative trajectory maps with those for ResNet50 (Figure 12) we find that the latter reveal a much more directional redundancy component to their trajectories as opposed to random walks. This further lends support to the thesis that optimization trajectories ensued when training neural networks are highly structured and have significant directional redudndancy.
Figure 12: Relative Trajectory Maps, with respect to initialization, of ResNet50 models for different amounts of momentum and weight decay.
An additional thing to note is that the above relative trajectory map for random walks covers the setting of decreasing the step size to mirror how the optimization procedure is setup for ResNet50. In the case of no such step size decay, the analytic and empirical versions of the relative trajectory map are depicted in the Figure 13.
Figure 13: No Step size decay: Relative Trajectory Map for a Random Walk/Brownian motion in both analytic and empirically simulated settings.
Appendix D Detailed Experimental Results
D.1 ResNet50: Switching off the hyperparameters
@@
The relative trajectory maps can be found in Figure 12.
@@
D.2 ResNet50: Weight Decay, AdamW
@@
Appendix E Directional Effects in Other Key Settings
In addition to the momentum and weight decay, there are other crucial hyperparameters, such as learning rate and batch size, whose directional effects warrant a mention. We carry out additional experiments in these settings, and from where, the key findings are that the learning rate, as it would be easy to guess, indeed encourages directional exploration leading to low MDS scores. But, somewhat more interestingly, we find that increasing the batch size also helps further exploration and thereby decreases the MDS scores. The trajectory maps can be found in the Figure 63 While we encourage the curious reader to have a look at the Appendix E.9, we find that with increased batch size, the angle between the updates as well as the angle between the update and the current location become increasingly obtuse, and thus making room for a wider directional exploration. In contrast, for smaller batch sizes these angles are closer to $90^{\circ}$. We hypothesize that a similar mutual interaction, as observed with weight decay and momentum, also occurs with batch size is considered. A detailed analysis, however, remains outside the current scope.
Lastly, we also experimented with Sharpness-Aware Minimization [Foret et al., 2021] (SAM), where we found that a higher value of the SAM regularization coefficient leads to a slightly increased directional similarity, which could potentially be related to SAM directing optimisation to flatter basins wherein the individual points are more directionally alike and have higher cosine similarities. The detailed results can be found in the Appendix E.3.
Other Settings and Datasets.
E.1 ViT: Weight Decay, AdamW
(a) $\angle(\theta_{t+1}-\theta_{t},\theta_{t})$
(b) $\angle(\theta_{t+1}-\theta_{t},\theta_{T}-\theta_{0})$
(c) $\angle(\theta_{t+k}-\theta_{t},\theta_{t}-\theta_{t-k})$, for $k=1$
(d) $\angle(\theta_{t}-\theta_{0},\theta_{T}-\theta_{0})$
(e) $\angle(\theta_{t+1}-\theta_{t},\theta_{T}-\theta_{0})$
(g) Apex Angle at Origin $\angle(\theta_{t},\theta_{0})$
(a) $\|\theta_{t}\|_{2}$
(b) $\|\theta_{t+k}-\theta_{t}\|_{2}$
(c) $\|\theta_{t}-\theta_{0}\|_{2}$
(a) Eigenvalues: ${\bf K}$
(b) Eigenvalues: ${\bf K}_{0}$
(c) Eigenvalues: ${\bf C}$
(d) Eigenvalues: ${\bf C}_{0}$
E.2 ResNet50: Weight Decay, SGD
@@
E.3 ResNet50: Sharpness Aware Minimization analysis
@@
E.4 ResNet50: Momentum Analysis, LR 0.1, WD 0.0001
@@
E.9 VGG16 Batch Size Analysis